OpenStudy (anonymous):

A car costs $25,750.00 and depreciates in value by 20% per year. How much will the car be worth after 5 years? $5,150.00 $8,437.76 $8,240.00 $20,600.00

OpenStudy (tkhunny):

After one year: 25750 * (1 - 0.20) = 25750 * 0.80 = 20600 After two years: 25750 * 0.80^2 = 16480 After three years: You tell me.

OpenStudy (anonymous):

C, right?

OpenStudy (tkhunny):

I don't know. I'm waiting for you to show your work and prove it to me.

OpenStudy (tkhunny):

I suspect, since we lose about a decimal place for each multiplication, that it is unlikely to end in "0.00".